Angela Raimondi wants to invest $8500,part at 14% annual interest and part at 12% annual interest.If she wants to earn the same amount of interest from eavh investment,how much should she invest at 14%?(Round the answer to the nearest cent.)
Answer by Nate(3500)

0.14(x) = 0.12(8500 - x)
0.14x = 1020 - 0.12x
0.26x = 1020
x = 3,923.08
8500 - 3923.08 = 4,576.92
$3,923.08 in 14%
$4,576.92 in 12%